# Efficient 3D Recognition with Event-driven Spike Sparse Convolution

### v2

@@ Introduction
Architecture Method Input $T\times D$ Param (M) Power (mJ) Accuracy (%) ANN PointNet (Qi et al. 2017)CVPR Point N/A 3.47 0.14 89.2 KPConv (Thomas et al. 2019)CVPR Point N/A 14.3 - 92.9 Pointformer (Zhao et al. 2021)ICCV Point N/A 4.91 0.11 93.7 3DShapeNets (Wu et al. 2015)CVPR Voxel N/A 6.92 0.15 77.3 3DVGG-B (Graham et al. 2017) Voxel N/A 5.23 0.12 88.2 E-3DSNN* Voxel N/A 3.27 0.17 90.9 SNN SpikePointNet (Ren et al. 2024)NeurIPS Point 4$\times$1 3.47 0.03 88.2 SpikingPointNet (Lan et al. 2023)ICCV Point 16$\times$1 3.47 0.13 88.6 P2SResLNet (Wu et al. 2024b)AAAI Point 4$\times$1 14.3 - 88.7 E-3DSNN (Ours) Voxel 1$\times$4 1.87 0.01 91.5 Voxel 1$\times$4 3.27 0.02 91.7
@@
As shown in Tab. 1, we compare our method with the previous state-of-the-art ANN and SNN domain. Notably, with only 3.27M parameters, the E-3DSNN achieves the best accuracy of 91.7%, regardless of voxel or point input in the SNN domain, showcasing significant advantages in both accuracy and efficiency. Specifically, E-3DSNN (This work) vs. SpikePointNet vs. SpikingPointNet: Power 0.01mJ vs. 0.03mJ vs. 0.13mJ; Param 1.87M vs. 3.47M vs. 3.47M; Accuracy 88.2% vs. 88.6% vs. 91.5%. That is, our model has +2.8% higher accuracy than SpikingPointNet (Lan et al. 2023) with only the previous 11.4% energy consumption. Moreover, the performance gap between SNNs and ANNs is significantly narrowed. For instance, under lower parameters, the performance of Pointformer (Zhao et al. 2021) and E-3DSNN are comparable, and the energy efficiency is 11$\times$.

### v3

@@ Introduction
Architecture Method Input $T\times D$ Param (M) Power (mJ) Accuracy (%) ANN PointNet (Qi et al. 2017)CVPR Point N/A 3.47 2.02 89.2 KPConv (Thomas et al. 2019)CVPR Point N/A 14.3 - 92.9 Pointformer (Zhao et al. 2021)ICCV Point N/A 4.91 5.1 93.7 3DShapeNets (Wu et al. 2015)CVPR Voxel N/A 6.92 0.61 77.3 3DVGG-B (Graham et al. 2017) Voxel N/A 5.23 0.45 88.2 E-3DSNN* Voxel N/A 3.27 0.30 90.9 SNN SpikePointNet (Ren et al. 2024)NeurIPS Point 4$\times$1 3.47 0.24 88.2 SpikingPointNet (Lan et al. 2023)ICCV Point 16$\times$1 3.47 0.91 88.6 P2SResLNet (Wu et al. 2024b)AAAI Point 4$\times$1 14.3 - 88.7 E-3DSNN (Ours) Voxel 1$\times$4 1.87 0.02 91.5 Voxel 1$\times$4 3.27 0.04 91.7
@@
As shown in Tab. 1, we compare our method with the previous state-of-the-art ANN and SNN domain. Notably, with only 3.27M parameters, the E-3DSNN achieves the best accuracy of 91.7%, regardless of voxel or point input in the SNN domain, showcasing significant advantages in both accuracy and efficiency. Specifically, E-3DSNN (This work) vs. SpikePointNet vs. SpikingPointNet: Power 0.02mJ vs. 0.24mJ vs. 0.91mJ; Param 1.87M vs. 3.47M vs. 3.47M; Accuracy 88.2% vs. 88.6% vs. 91.5%. That is, our model has +2.8% higher accuracy than SpikingPointNet (Lan et al. 2023) with only the previous 15.4% energy consumption. Moreover, the performance gap between SNNs and ANNs is significantly narrowed. For instance, under lower parameters, the performance of Pointformer (Zhao et al. 2021) and E-3DSNN are comparable, and the energy efficiency is 127$\times$.